OpenStudy (anonymous):

Two blocks sit on a level, frictionless surface. Block A has a mass of 0.5 kg and block B has a mass of 2.5 kg. The blocks are forced together, compressing a massless spring between them. The system is then released from rest. The spring is not fastened to either block, so it drops to the surface after it has expanded. Block A acquires a speed of 3 m/s. How much potential energy was stored initially in the compressed spring?

OpenStudy (anonymous):

The two msses must have the same kinectic energy, because the spring would b=make the same force on both, but the velocities will be different, because of the masses. Since all the energy contained in the spring in the begining was the potential energy, you just need to find the total energy after the expansion, because energy conserves.

OpenStudy (anonymous):

so by using KE=1/2mv^2 I found that KE=2.25 => speed of Block B=> 2.25 = 1/2(2.5)v^2 => v=1.342 what do I do after that??

OpenStudy (anonymous):

You don't need to find the speed, since the kinectic energy must be the same for both, to get the total energy you just need to multiplicate the kinectic energy you found by 2 and that should already give you the potential energy.

OpenStudy (anonymous):

the answer is USpring = 2.7 J. can you explain the answer for me please

OpenStudy (anonymous):

I'm thinking, did you got to this number or it is the given answer?

OpenStudy (anonymous):

its the given answer.

OpenStudy (anonymous):

hahaha then I have to think, I might be wrong about the kinectic energies being the same. Just a minute.

OpenStudy (anonymous):

ok thank you :)

OpenStudy (anonymous):

Ok I think I have something, since F=dp/dt, where p is the momentum, what is the same for both is the momentum, not the kinectic energy, knowing the momentum of one, you find the velocity of the other, calculate both kinectic energies and sum them to get the total energy, sich is also the potential at the begining.

OpenStudy (anonymous):

thank you very much :) I really appreciate it :)